Daily high temperatures in St. Louis for the last week were as​ follows: 95​, 94​, 95​, 95​, 95​, 90​, 90 ​(yesterday).

​a) The high temperature for today using a​ 3-day moving average ​= nothing degrees ​(round your response to one decimal​ place).

​b) The high temperature for today using a​ 2-day moving average ​= nothing degrees ​(round your response to one decimal​ place).

​c) The mean absolute deviation based on a​ 2-day moving average​ = nothing degrees ​(round your response to one decimal​ place). ​

d) The mean squared error for the​ 2-day moving average​ = nothing degrees squared ​(round your response to one decimal​ place). ​

e) The mean absolute percent error​ (MAPE) for the​ 2-day moving average​ = nothing​% ​(round your response to one decimal place​).
Answer #1

Temperature

3-Day MA

2-Day MA

MAD

MSE

MAPE

95

94

=(95+94)/2

=ABS((94-94.5))

=(94-93.43)^2

=(ABS((94-93.43)/94))*100

95

=(95+94+95)/3

=(94+95)/2

=ABS((95-94.5))

=(95-93.43)^2

=(ABS((95-93.43)/95))*100

95

=(94+95+95)/3

=(95+95)/2

=ABS((95-95))

=(95-93.43)^2

=(ABS((95-93.43)/95))*100

95

=(95+95+95)/3

=(95+95)/2

=ABS((95-95))

=(95-93.43)^2

=(ABS((95-93.43)/95))*100

90

=(95+95+90)/3

=(95+90)/2

=ABS((90-92.5))

=(90-93.43)^2

=(ABS((90-93.43)/90))*100

90

=(95+90+90)/3

=(90+90)/2

=ABS((90-91.67))

=(90-93.43)^2

=(ABS((90-93.43)/90))*100

Average

=SUM(95+94+95+95+95+90+90)/7

= 93.4

=SUM(91.67+93.33+95)/3

= 93.3

=SUM(94.5+94.5+95+95+92.5+90)/6

= 93.6

=SUM(0.5+0.5+0+0+2.5+1.67)/6 =0.9

=SUM(0.32+2.46+2.46+2.46+11.76+11.76)/6 =5.2

=(SUM(0.61+1.65+1.65+1.65+3.81+3.81)/6)

= 2.2

a) SUM(91.67+93.33+95)/3 = 93.4

b) SUM(94.5+94.5+95+95+92.5+90)/6 = 93.6

c) SUM(0.5+0.5+0+0+2.5+1.67)/6 = 0.9

d) SUM(0.32+2.46+2.46+2.46+11.76+11.76)/6 = 5.2

e) SUM(0.61+1.65+1.65+1.65+3.81+3.81)/6) = 2.2